The force between two charges is 3.5 x 10^»N. What is the force if the distance separating the charges is doubled. (4pts)
Answer #1

We know that the Force between the two charge

F=kQq/r^2. ...1

According to the question R =2r

Put in above equation

F' =kQq/(2r)^2

F' = kQq/4r^2 ...2

From equation 1 and 2

F' =F/4

here F=3.5×10^(-2) ( first varify red digit it is not so clear )

So

F' =3.5×10^(-2)/4

F' = 0.875×10^(-2) N. Answer.